### v1

# Fine-grained Semantics Integration for Large Language Model-based Recommendation

Abstract.
Recent advances in Large Language Models (LLMs) have shifted in recommendation systems from the discriminative paradigm to the LLM-based generative paradigm, where the recommender autoregressively generates sequences of semantic identifiers (SIDs) for target items conditioned on historical interaction. While prevalent LLM-based recommenders have demonstrated performance gains by aligning pretrained LLMs between the language space and the SID space, modeling the SID space still faces two fundamental challenges: (i) Semantically Meaningless Initialization: SID tokens are randomly initialized, severing the semantic linkage between the SID space and the pretrained language space at start point, and (ii) Coarse-grained Alignment: existing SFT-based alignment tasks primarily focus on item-level optimization, while overlooking the semantics of individual tokens within SID sequences.
To address these challenges, we propose TS-Rec, which can integrate Token-level Semantics into LLM-based Recommenders. Specifically, TS-Rec comprises two key components: (1) Semantic-Aware embedding Initialization (SA-Init), which initializes SID token embeddings by applying mean pooling to the pretrained embeddings of keywords extracted by a teacher model; and (2) Token-level Semantic Alignment (TS-Align), which aligns individual tokens within the SID sequence with the shared semantics of the corresponding item clusters. Extensive experiments on two real-world benchmarks demonstrate that TS-Rec consistently outperforms traditional and generative baselines across all standard metrics. The results demonstrate that integrating fine-grained semantic information significantly enhances the performance of LLM-based generative recommenders.
00footnotetext: Corresponding Author.
Language Models, Generative Recommendation, Semantic IDs
$c$s: Information systems Recommender systems
Figure 1. Representations of SIDs before (i.e., left column) and after (i.e., right column) alignment. The randomly initialized SIDs are initially far away from the semantic space of the LLM (i.e., grey points), while SA-Init enables SIDs within the semantic space of the LLM. This makes the SID alignment of the randomly initialized one more difficult than SA-Init.
1. Introduction
Large language model (LLM)-based generative recommendation has recently demonstrated great potential in modeling user preferences (Zheng et al., 2024; Bao et al., 2023; Liao et al., 2024; Sheng et al., 2024a; Wu et al., 2024). In generative recommendation, items are represented as discrete tokens rather than continuous vectors(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023; Zhai et al., 2024), enabling the model to directly generate the token corresponding to the next item through an autoregressive process. Building on this paradigm, LLM-based generative recommendation further incorporates large language models to leverage their rich world knowledge, enhancing the recommender’s understanding of items and thereby improving recommendation performance(Liu et al., 2025; Han et al., 2025; Team, 2025; Zhang et al., 2023).
In a typical LLM-based generative recommendation pipeline (Liu et al., 2025; Kong et al., 2025b; You et al., 2025), the training procedure can be broadly divided into three stages: (1) item tokenization as semantic identifiers (SIDs)(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023): each item is first represented as a sequence of $N$ fixed-length tokens (i.e., SIDs) that lie outside the original vocabulary of the LLM (e.g., three tokens like ¡a_236¿ ¡b_231¿ ¡c_226¿), commonly obtained using vector quantization methods such as RQ-VAE(Zeghidour et al., 2021); (2) alignment(Zheng et al., 2024; Wang et al., 2024) between these out-of-vocabulary (OOV) item tokens and the LLM backbone through supervised fine-tuning (SFT), enabling the model to understand these tokens for recommendation; and (3) preference learning (e.g., reinforcement learning (RL)(Tan et al., 2025; Kong et al., 2025b; Team, 2025; Zhou et al., 2025; Kong et al., 2025a; You et al., 2025) or direct preference optimization (DPO) (Chen et al., 2024; Bai et al., 2024; Rafailov et al., 2023; Ye et al., 2025)) to further enhance the user preferences modeling capabilities, thereby pushing the upper bound of recommendation quality. As the first two steps, item tokenization and alignment with LLMs directly determine how the LLM understands items and thus largely affect recommendation performance, playing vital roles in LLM-based generative recommendation. However, existing approaches (Rajput et al., 2023; Zheng et al., 2024; Kong et al., 2025b) for these two steps remain coarse, which limits the LLM’s ability to effectively understand OOV item tokens and consequently leads to suboptimal recommendation performance. Specifically, we identify two key limitations in current item tokenization and alignment with LLMs as follows:
  • Semantically Meaningless Initialization. In the item tokenization stage, SIDs are typically initialized in a semantically meaningless manner, deviating from the LLM representation space and hindering their subsequent alignment with LLMs. As illustrated in Figure 1(a), the current SID initialization strategy relies on random initialization, resulting in representations that deviate substantially from the intrinsic representation space of the LLM. Consequently, aligning these SIDs to the LLM becomes more challenging, as they must be pulled from a distant region of the representation space (cf. Figure 1(a)).
  • Coarse-grained Alignment. In alignment with LLM stage, current alignment methods tend to adopt a coarse-grained manner (Rajput et al., 2023; Zheng et al., 2024; Kong et al., 2025b), where the alignment task is constructed to translate the SID into natural language (e.g., translate ¡a_236¿ ¡b_231¿ ¡c_226¿ into “Relay/Capacitor Hard Start Kit”). However, this alignment strategy is relatively coarse and lacks a fine-grained understanding of SIDs at the token level. As illustrated in Figure 2, when we ask the LLM what general feature the first token ¡a_236¿ represents, it instead responds with a specific item, which is unsatisfactory. In practice, however, a token at the SID level typically corresponds to a broader category or attribute, rather than an individual item. This mismatch indicates that the LLM fails to capture the fine-grained semantics at the token level, resulting in insufficient token-level understanding and further hindering the recommendation performance.
Figure 2. TS-Rec enables recommenders to have high-level and fine-grained understanding of token-level SIDs.
We argue that the fine-grained semantics integration is crucial in both item tokenization and alignment with LLM stages, while current approaches largely ignore such token-level semantics. To address the above limitations, we introduce TS-Rec, a generative recommendation framework that enhances SID comprehension, which comprises two key components:
  • Semantic-Aware Embedding Initialization. We initialize SID token embeddings in a semantic-aware manner by mean pooling the pretrained embeddings of keyword sequence extracted by a semantic extractor. Specifically, for each token, the semantic extractor summarizes the common properties shared among all items containing that token, generating a corresponding token description and keywords. We then initialize the SID token embeddings via the mean pooling of the pre-trained embeddings of keyword sequences. This ensures that each SID token is endowed with meaningful semantics prior to post-alignment, thus providing a better starting point for the subsequent process.
  • Token-level Semantic Alignment. Building upon standard SFT objectives, we formulate a token-level alignment task to align the SID space and the language space at a finer granularity. Specifically, we use the token descriptions generated by the semantic extractor and their corresponding SID tokens as training pairs. This alignment objective ensures that the model captures both the semantics of individual SID tokens and complete SID sequences. By grounding these token-wise semantic units in the LLM’s language space, the alignment strengthens the LLM’s fine-grained understanding of SIDs and promotes their effective use in generation process.
To evaluate the performance of TS-Rec, we conduct extensive experiments on two widely-adopted real-world benchmarks, Amazon Industrial and Office (Ni et al., 2019). Empirical results demonstrate that TS-Rec consistently achieves superior performance across all standard metrics, yielding significant relative improvements of up to 7.17% and 4.27% on the two datasets, respectively. Moreover, our analysis reveals that the superiority of TS-Rec stems from its ability to bridge the semantic gap between discrete semantic identifiers and the LLM’s language space, providing a more robust optimization starting point and a more granular understanding of SIDs. Furthermore, we show that TS-Rec serves as a superior semantic backbone that significantly elevates the performance ceiling of the subsequent ”SFT-then-RL” pipeline (Kong et al., 2025b). These findings underscore the critical importance of fine-grained semantic integration in LLMs in recommendation task.
2. Related Work
Traditional recommendation systems have been dominated by discriminative paradigms, such as collaborative filtering (He et al., 2017; Sarwar et al., 2001) and click-through rate (CTR) models (Zhou et al., 2018, 2019; Wu et al., 2025; Sheng et al., 2024b). Recently, the field has experienced a paradigm shift toward generative recommendation (Zhang et al., 2025; Rajput et al., 2023; Zhai et al., 2024; Zhou et al., 2025; Zheng et al., 2024; Team, 2025; Hao et al., 2025; Han et al., 2025), which reframes recommendation as a next-item generation problem, where transformer-based models (Yang et al., 2024; Raffel et al., 2020) autoregressively generate recommended items conditioned on user interaction histories. This paradigm offers greater generalizability and significantly improves performance in cold-start and long-tail scenarios.
2.1. SID-based Generative Recommendation
A cornerstone of generative paradigm is the design of Semantic Identifiers (SIDs), which serve as discrete tokens bridging continuous item multimodal representations and discrete item identifiers.
Early pioneer work TIGER (Rajput et al., 2023) utilizes residual quantization (RQ-VAE) (Lee et al., 2022) to derive hierarchical SIDs from item metadata. Building upon this foundation, more effective codebook construction strategies are explored to obtain higher-quality SIDs (Luo et al., 2025; Wang et al., 2025b; Chen et al., 2025). Subsequent studies including LETTER (Wang et al., 2024), EAGER (Wang et al., 2025a), and UNGER (Xiao et al., 2025) has further integrated collaborative signals into the generative process. In industrial settings, HSTU (Zhai et al., 2024) optimizes the sequence modeling efficiency through a hierarchical transformer architecture tailored for high-throughput recommendation. OneRec (Zhou et al., 2025) further establishes an end-to-end training pipeline to effectively deploy generative recommendation in large-scale industrial systems.
Despite these advancements, traditional SID-based generative methods only encode item semantics into the random initialized SID embeddings, while largely overlooking the potential of exploiting the inherent semantic capabilities of the generative backbone.
2.2. LLM-based Generative Recommendation
Another line of work focuses on utilizing modern LLMs (Yang et al., 2024; Team, 2024) as the recommendation backbone to harness their vast semantic understanding and world knowledge (Bao et al., 2023; Liao et al., 2024; Karra and Tulabandhula, 2024; Liu et al., 2025; Team, 2025; Kong et al., 2024).
Early efforts like TallRec (Bao et al., 2023) and LLaRa (Liao et al., 2024) demonstrate the effectiveness of instruction tuning, employing item titles to align LLMs with recommendation objectives via specialized prompts. However, relying solely on textual titles to represent items has notable limitations: Long or blurry titles often strain the LLM’s context window and exacerbate the risk of hallucinations. To address this, LC-Rec (Zheng et al., 2024) and OneRec-Think (Liu et al., 2025; Team, 2025) adpot compact SID-based representations for items, and align the representation spaces of pre-trained LLMs and SIDs through token expansion and multi-task supervised fine-tuning. Building on these SFT-aligned models, reinforcement learning (RL) algorithms like GRPO (Shao et al., 2024) are further introduced to refine recommendation quality (Team, 2025; Kong et al., 2025b; Tan et al., 2025).
Nevertheless, existing approaches typically initialize newly added SID tokens randomly and align LLMs only at the level of complete hierarchical SIDs, overlooking fine-grained semantic learning at the individual token level. Since RL optimization relies on the SFT model as its starting point, insufficient SID understanding at this stage can limit subsequent gains. These limitations hinder effective SID comprehension and motivate our work.
Figure 3. The overall framework of our TS-Rec. The pipeline consists of two complementary modules: (1) SA-Init, which initializes SID embeddings with semantic descriptions generated by an Semantic Extractor from clustered items; (2) TS-Align, which aligns each SID token with its semantics through multi-task instruction tuning.
3. Preliminaries
This section introduces key concepts and notations used throughout the paper, focusing on the formulation of generative recommendation, inculuding the tokenization mechanism and the SFT objective.
3.1. Task Formulation
Let $H_{u}$ be the chronologically ordered interaction history of user $u$, Each item $i\in H_{u}$ is represented as a three-level SID tuple $e_{i}=\{s^{(1)}_{i},s^{(2)}_{i},s^{(3)}_{i}\}$, where $s^{(l)}_{i}$ is the SID token at level $l$. Given $H_{u}$, the generative recommender $\pi_{\theta}$, parameterized by $\theta$, is trained to predict the preferred item $i_{p}$ that matches the preference of user $u$ from the item set.
3.2. Tokenization and Embedding Initialization
3.2.1. Item Tokenization.
For each item $i$, we concatenate its textual title and description into a single sentence, which is then fed into a frozen encoder to produce a $d$-dimensional embedding $\mathbf{x}\in\mathbb{R}^{d}$. The continuous embedding is subsequently discretized using the RQ-KMeans algorithm (Luo et al., 2025).
Let $\tilde{\mathbf{M}}=[\mathbf{x}_{1};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ be the matrix stacking the embeddings of all $N$ items. We initialize $\mathbf{R}^{(1)}=\tilde{\mathbf{M}}$. For each layer $l\in\{1,\dots,L\}$, where $L$ is the number of hierarchical levels in the semantic codebook, we learn a codebook $\mathbf{C}^{(l)}=\{\mathbf{c}_{k}^{(l)}\}_{k=1}^{K_{l}}$ by running K-means with $K_{l}$ centroids on the current residuals $\mathbf{R}^{(l)}$:
(1) $\mathbf{C}^{(l)}=\text{K-means}(\mathbf{R}^{(l)},K_{l}).$
For item $i$ ($1\leq i\leq N$), the the nearest centroid index is assigned as:
(2) $s_{i}^{(l)}=\arg\min_{k}\|\mathbf{R}_{i}^{(l)}-\mathbf{c}_{k}^{(l)}\|_{2},$
where $\|\cdot\|$ denotes the Euclidean norm.The residual is updated as:
(3) $\mathbf{R}_{i}^{(l+1)}=\mathbf{R}_{i}^{(l)}-\mathbf{c}_{s_{i}^{(l)}}^{(l)}.$
We set the number of codebook levels to $L=3$. Each item is thus represented by a sequence of semantic identifiers $\{s^{(1)}_{i},s^{(2)}_{i},s^{(3)}_{i}\}$, which serves as a unique [^1] To avoid collisions, we reassign the last SID token randomly for any conflicting items. token sequence for each item $i$ and will be used by the recommender model in autoregressive generation.
3.2.2. Embedding Initialization.
Following the tokenization process, the newly generated SID tokens in $\mathcal{V}_{\text{SID}}$ are integrated into the pre-trained vocabulary of the LLM, denoted as $\mathcal{V}_{\text{Pre}}$, to form an expanded vocabulary:
(4) $\mathcal{V}=\mathcal{V}_{\text{SID}}\cup\mathcal{V}_{\text{Pre}}.$
An embedding vector is then assigned to each newly added token in $\mathcal{V}_{\text{SID}}$. These embeddings are typically initialized by sampling from a Gaussian distribution $\mathcal{N}(\mu,\Sigma)$ fitted to the pre-trained token embeddings. Here, $\mu$ and $\Sigma$ denote the mean vector and covariance matrix of the LLM’s pre-trained embedding space.
3.3. Supervised Fine Tuning for LLM-based Recommenders
In supervised fine-tuning (SFT), the semantic understanding and generative capabilities of large language models (LLMs) are leveraged to enhance sequential recommendation. As detailed in Appendix A.2, the fine tuning tasks can be conveniently formulated as conditional language generation in a sequence-to-sequence manner. LLM-based recommenders are fine-tuned using the following language modeling objective:
(5) $\mathcal{L}_{\text{SFT}}=-\sum_{(X,Y)}\sum^{|Y|}_{j=1}\log\pi_{\theta}(Y_{j}\mid X,Y_{<j}),$
where $(X,Y)$ denotes a pair of instruction and target response from the training data, $Y_{j}$ denotes the $j$-th token in the response $Y$, and $Y_{<j}$ denotes the prefix tokens preceding $Y_{j}$. The recommender $\pi_{\theta}$ models the conditional probability of generating the target response $Y$ given instruction $X$.
4. Method
As illustrated in Figure 3, we propose a framework to bridge the semantic gap between SIDs and pre-trained LLMs. The method comprises two complementary modules: 1) SA-Init (Sec. 4.1), which statically injects semantic priors into SIDs using an Extractor LLM to address the cold-start issue; and 2) TS-Align (Sec. 4.2), which dynamically refines these representations through multi-task instruction tuning.
4.1. SA-Init: Semantic-Aware embedding Initialization
As discussed in Section 3.2, unlike natural language tokens, the SID tokens in $\mathcal{V}_{\text{SID}}$ are randomly initialized and lack inherent semantic meaning at the onset of training. This semantic gap often leads to suboptimal convergence. To bridge this gap, we introduce Semantic-Aware Embedding Initialization (SA-Init). By employing an Extractor LLM $\mathcal{M}_{ext}$ to synthesize the common semantics associated with each SID token, SA-Init injects semantic priors into the SID tokens, ensuring they are semantically grounded before the fine-tuning stage. Specifically, the SA-Init comprises the following steps:
  • Token-Specific Item Clusters. We first derive the SID vocabulary $\mathcal{V}_{\text{SID}}$ from the global item set $\mathcal{I}$. For each unique SID token $s\in\mathcal{V}_{\text{SID}}$, we aggregate the set of items that contain this token at any level of their identifier hierarchy. Formally, the cluster for token $s$ is defined as:
    (6) $\mathcal{I}_{s}=\{i\in\mathcal{I}\mid s\in e_{i}\},$
    where $e_{i}$ represents the SID tuple of item $i$. This step groups items based on shared structural codes, serving as a proxy for shared semantics.
  • Semantic Extraction. To interpret the semantic rationale behind the structural cluster $\mathcal{I}_{s}$, we first randomly sample a subset of items $\hat{\mathcal{I}}_{s}$. We then prompt the Extractor LLM $\mathcal{M}_{\text{ext}}$ (e.g., DeepSeek) to generate a textual description $\mathcal{D}_{s}$ capturing the shared characteristics, along with a concise list of keywords $\mathcal{W}_{s}$:
    (7) $\mathcal{D}_{s},\mathcal{W}_{s}=\mathcal{M}_{\text{ext}}(\mathcal{P}_{\text{ext}}(\hat{\mathcal{I}}_{s})),$
    where $\mathcal{P}_{\text{ext}}$ denotes the prompt template for semantic extraction (see Appendix A.1 for details).
  • Initialization via Keyword Aggregation. Finally, we bridge the gap between the extracted symbolic keywords and the continuous embedding space of the recommender. We concatenate the keywords in $\mathcal{W}_{s}$ and project them into the backbone model’s vocabulary space via tokenization, yielding a sub-token sequence $\mathcal{T}_{s}$:
    (8) $\mathcal{T}_{s}=\text{Tok}_{\text{backbone}}(\text{Concat}(\mathcal{W}_{s})),$
    where $\mathcal{T}_{s}$ consists of token IDs from the pre-trained vocabulary. To endow the SID token $s$ with an informative initial state, we perform an embedding lookup for each sub-token $v\in\mathcal{T}_{s}$ in the pre-trained embedding matrix $\mathbf{E}$. The embedding of $s$ is initialized by aggregating these semantic priors:
    (9) $\mathbf{e}_{s}=\frac{1}{|\mathcal{T}_{s}|}\sum_{v\in\mathcal{T}_{s}}\mathbf{E}[v].$
    This aggregation effectively aligns the initialized SID embeddings with the well-structured semantic space of the pre-trained LLM.
4.2. TS-Align: Token-level Semantic Alignment
To equip the LLM with a fine-grained understanding of the SID structure, we introduce Token-level Semantic Alignment (TS-Align). Standard alignment approaches typically align the entire SID sequence of an item with its global description (e.g., title). However, this holistic alignment often overlooks the specific semantic contribution of individual tokens. In contrast, TS-Align aligns each individual SID token $s$ with its distilled semantic description $\mathcal{D}_{s}$ (derived in Section 4.1). This establishes an explicit mapping between the atomic structural units and their semantic meanings.
Specifically, we construct a token-level instruction tuning dataset $\mathcal{D}_{\text{token}}$, consisting of bidirectional tasks formatted as instruction-response pairs $(x,y)$:
Task 1: Semantic-to-Token Alignment (Understanding) Instruction ($x$): Identify the specific SID token shared by items that exhibit the following scope and characteristics: $\mathcal{D}_{s}$ Response ($y$): $s$
Task 2: Token-to-Semantic Alignment (Generation) Instruction ($x$): Describe the typical scope and shared features of items associated with the token: $s$ Response ($y$): $\mathcal{D}_{s}$
Multi-Task Joint Optimization. Following the multi-task learning paradigm established in generative recommendation (xxx), we treat the sequential recommendation as the primary task and the proposed TS-Align tasks as auxiliary sub-tasks. During the SFT phase, we jointly optimize the model on both the recommendation data $\mathcal{D}_{\text{rec}}$ and the token alignment data $\mathcal{D}_{\text{token}}$. By integrating these auxiliary alignment objectives, the model is regularized to maintain semantic consistency within the SID representations while minimizing the standard next-token prediction loss. This joint training ensures that the learned SIDs are not only distinct identifiers for recommendation but also semantically grounded representations.
5. Experiments
In this section, we first describe our experimental setup, then present the results and analyze our approach to answer the following research questions:
  • RQ1: How does TS-Rec perform compared to other baselines?
  • RQ2: How do the key components of TS-Rec contribute to its recommendation performance?
  • RQ3: Does integrating reinforcement learning (RL)–based methods into TS-Rec lead to greater performance improvements?
  • RQ4: Can TS-Rec enhance the LLM’s understanding of SIDs?
5.1. Experimental Setting
Table 1. Statistics of the preprocessed datasets.
Datasets #Users #Items #Interactions Sparsity
Industrial 7,694 3,685 45,324 99.84%
Office 8,328 3,459 48,656 99.83%
5.1.1. Datasets.
We evaluate the performance of our proposed approach on two public subsets of the Amazon review dataset (Ni et al., 2019), specifically ”Industrial and Scientific” and ”Office Products”. Each item is associated with a title and a description. Following prior work(Zheng et al., 2024), we filter out users and items with fewer than five interactions. User interaction sequences are then constructed in chronological order. The statistics of the preprocessed datasets are summarized in Table 1.
5.1.2. Evaluation Settings.
To evaluate performance, we adopt two widely-used metrics including top-$K$ Hit Ratio (HR@$K$) and top-$K$ Normalized Discounted Cumulative Gain (NDCG@K@K), with $K$ set to 3, 5, and 10. We use a full ranking evaluation on the entire item set rather than a sample-based evaluation. For generative baselines that rely on beam search, the beam size is uniformly fixed at 20.
5.1.3. Baselines.
We compare proposed approach against two categories of leading recommendation baselines. These baselines can be broadly categorized into: (1) Traditional Recommenders: including GRU4Rec, Caser and SASRec. (2) Generative Recommenders: including TIGER, HSTU and LC-Rec.
The traditional recommendation models are as follows:
  • GRU4Rec (Hidasi and Karatzoglou, 2018) is a CNN-based model designed to capture both long-term and short-term user interests.
  • Caser (Tang and Wang, 2018) is an RNN-based recommendation model that is relatively simple yet effective.
  • SASRec (Kang and McAuley, 2018) is a widely used sequential recommendation model implemented with self-attention, enabling it to capture dynamic user preferences over time.
The generative recommendation models are as follows:
  • TIGER (Rajput et al., 2023) adopts a generative retrieval paradigm for recommendation and introduces Semantic IDs to index items.
  • HSTU (Zhai et al., 2024) employs a streaming architecture tailored for high-cardinality and non-stationary recommendation data.
  • LC-Rec (Zheng et al., 2024) is an LLM-based recommendation model that incorporates a series of tasks to align the SID space with the language space, thereby leveraging world knowledge to boost recommendation performance.
Table 2. Recommendation performance on two real-world datasets. The best performance is highlighted in boldface, while the second-best performance is underlined.
Dataset Models HR@3 NDCG@3 HR@5 NDCG@5 HR@10 NDCG@10
Industrial GRU4Rec 0.0638 0.0542 0.0774 0.0598 0.0999 0.0669
Caser 0.0618 0.0514 0.0717 0.0555 0.0942 0.0628
SASRec 0.0790 0.0700 0.0909 0.0748 0.1088 0.0806
TIGER 0.0852 0.0742 0.1010 0.0807 0.1321 0.0908
HSTU 0.0927 0.0885 0.1037 0.0918 0.1163 0.0958
LC-Rec 0.0915 0.0805 0.1057 0.0862 0.1332 0.0952
TS-Rec 0.0966 0.0853 0.1153 0.0930 0.1462 0.1029
Office GRU4Rec 0.0629 0.0528 0.0789 0.0595 0.1019 0.0669
Caser 0.0748 0.0615 0.0865 0.0664 0.1093 0.0737
SASRec 0.0861 0.0769 0.0949 0.0805 0.1120 0.0858
TIGER 0.0986 0.0852 0.1163 0.0960 0.1408 0.1002
HSTU 0.1134 0.1031 0.1252 0.1079 0.1400 0.1126
LC-Rec 0.0921 0.0807 0.1048 0.0859 0.1237 0.0920
TS-Rec 0.1200 0.1056 0.1307 0.1100 0.1541 0.1175
5.1.4. Implementation Details.
For constructing the item indices, we utilize the Qwen3-Embedding-4B encoder to generate continuous embeddings for item titles and descriptions. The codebook depth $L$ is set to 3, where each layer contains 256 codebook vectors with a dimensionality 32, optimizaed via the RQ-KMeans algorithm. Notably, to mitigate indices collision, we reassign the last SID token randomly for any conflicting items.
For the LLM finetuning, we implement our proposed approach using Qwen2.5-1.5B (Yang et al., 2024). The learning rate is set to 3e-4 with a total batch size of 1024. We train for 3 epochs on each dataset. To prevent overfitting, we apply an early stopping strategy that terminates training if the loss on the validation set does not decrease for three consecutive evaluation steps. All experiments are conducted using 8 NVIDIA H20 96GB GPUs.
5.2. Overall Performance (RQ1)
The experimental results of applying TS-Rec to Qwen2.5‑1.5B across two real‑world benchmarks are presented in Table 2. We summarize our key observations as follows:
(1) TS-Rec consistently achieves superior performance across both benchmarks, outperforming all compared baselines, including both traditional and generative recommendation models. On Amazon Industrial, TS-Rec achieves relative gains of +7.17% in HitRatio and NDCG, while on Amazon Office the relative gains reach 4.27%. Beyond these numerical improvements, we attribute the superior performance to two core factors: (a) Better Initialization: SA‑Init provides the LLM with a strong semantic grounding at the token‑embedding level, enabling more effective optimization in subsequent training. (b) Fine‑grained Semantic Alignment: TS‑Align ensures that nuanced SID‑level semantic relations are preserved and leveraged during the recommendation process, yielding more accurate ranking predictions. These results validate the integration of fine‑grained semantic knowledge as a key driver for enhancing LLM‑based recommenders.
(2) LLM‑based recommendation models consistently outperform traditional baselines across datasets, reflecting the advantages brought by natural language reasoning capabilities and rich world knowledge encoded in large language models. Nevertheless, effectively adapting LLMs to recommendation tasks remains non‑trivial. Challenges include aligning general‑purpose language knowledge with domain‑specific recommendation objectives, and balancing generation flexibility with ranking precision. Our findings suggest that TS-Rec, by providing a stronger starting point and explicit semantic integration, offers a promising pathway to address these challenges.
Figure 4. Training dynamics of TS-Rec with different SA-Init depths on the Industrial and Office datasets. Configurations include Random Init, SA-Init 1 (first layer), SA-Init 2 (first two layers), and SA-Init 3 (all three layers).
5.3. Ablation Study (RQ2)
In this section, we conduct ablation studies to validate the contribution of the core components in TS-Rec. Our analysis focuses on two key aspects: (1) assessing the individual and combined impact of SA‑Init and TS‑Align, and (2) analyzing the influence of token initialization depth on the overall recommendation performance.
5.3.1. Impact of SA-Init and TS-Align.
To assess the contribution of each component in TS-Rec as well as their combined effect, we conduct a comprehensive ablation study. The results are summarized in Table 3, and the key findings are as follows: (1) Both SA‑Init and TS‑Align contribute substantively to performance gains. Removing SA‑Init results in notable performance degradation. On Industrial, HR@5 drops by 1.90% (from 0.1153 to 0.1131) and NDCG@5 by 2.04% (from 0.0930 to 0.0911). On Office, the drops are more pronounced: HR@5 decreases by 8.33% (from 0.1307 to 0.1198) and NDCG@5 by 13.36% (from 0.1100 to 0.0950). Similarly, removing TS‑Align leads to performance declines. On Industrial, HR@5 falls by 3.64% (from 0.1153 to 0.1111) and NDCG@5 by 0.21% (from 0.0930 to 0.0928), while on Office, HR@5 decreases by 0.53% (from 0.1307 to 0.1300) and NDCG@5 by 1.45% (from 0.1100 to 0.1084). These results indicate that: (a) SA‑Init provides a strong initialization for token embeddings, facilitating more effective optimization during post-training. (b) TS‑Align injects fine-grained semantic coherence into SID representation, boosting downstream ranking accuracy. (2) The combination of SA-Init and TS-Align yields the strongest SFT-based optimization. This synergy suggests that SA-Init and TS-Align are complementary components, as each addresses different aspects of model optimization, and together they drive maximal performance gains.
Table 3. Ablation studies on the Industrial and Office datasets. ”HR@5” is an abbreviation for HitRatio@5. The best performance is highlighted in boldface.
Method Industrial Office
HR@5 NDCG@5 HR@5 NDCG@5
TS-Rec 0.1153 0.0930 0.1307 0.1100
w/o SA-Init 0.1131 0.0911 0.1198 0.0950
w/o TS-Align 0.1111 0.0928 0.1300 0.1084
w/o SA-Init & TS-Align 0.1080 0.0860 0.1278 0.1045
5.3.2. Impact of token initialization Depth.
Although the preceding results demonstrate the effectiveness of SA-Init, the underlying mechanisms through which it influences training remain unclear. We therefore conduct a set of controlled experiments to investigate the aspects of training affected by SA-Init. Specifically, we compare the following four initialization strategies for the SID token embeddings:
  • Random Init: Initialize SID token embeddings using the Qwen2.5’s default initialization method.
  • SA-Init 1: Apply SA-Init to initialize the first layer of the SID sequence’s token embeddings.
  • SA-Init 2: Apply SA-Init to initialize the first two layers of the SID sequence’s token embeddings.
  • SA-Init 3: Apply SA-Init to initialize all three layers of the SID sequence’s token embeddings.
We track training dynamics on both the Industrial and Office datasets, recording the eval loss and recommendation performance in terms of HR@5 throughout the training process. As illustrated in Figure 4, our main obserbvations are: (1) Deeper SA‑Init reduces early-stage loss. On both datasets, models initialized with SA‑Init exhibit substantially lower eval loss from the outset compared to Random Init, with SA‑Init 3 converging to the lowest final loss values. On Industrial, early-phase loss drops by approximately 0.5 to 1.0 points, while Office shows similar magnitudes of reduction. (2) Faster and higher recommendation performance convergence. HR@5 improves more quickly and to higher levels as initialization depth increases. On Office, SA‑Init 3 surpasses 0.12 HR@5 before step 100, while Random Init only approaches this level at the end of training. On Industrial, SA‑Init 3 consistently maintains a gap over SA‑Init 2 and SA‑Init 1 across all steps. (3) Monotonic performance scaling with depth. Across both datasets, the performance ordering is consistent: SA‑Init 3 ¿ SA‑Init 2 ¿ SA‑Init 1 ¿ Random Init. This aligns with the expectation that deeper semantic grounding in early layers provides sustained optimization advantages. (4) Cross-domain generalization. Despite domain differences (Industrial vs. Office), trends are robust—indicating that the benefits of deeper SA‑Init are generalizable to recommendation environments with heterogeneous item spaces.
5.4. Integration of Reinforcement Learning (RQ3)
Recent breakthroughs in Large Language Models, exemplified by the DeepSeek (Guo et al., 2025) series, have sparked a significant trend in the recommendation community toward adopting the SFT-then-RL paradigm (Kong et al., 2025b). In this workflow, the Supervised Fine-Tuning (SFT) stage establishes the foundational policy, while Reinforcement Learning (RL) further aligns this policy with user preferences and ranking metrics. However, RL optimization is notoriously sensitive to policy initialization; a backbone that lacks a deep understanding of the semantic identifiers space often leads to unstable reward signals and suboptimal convergence.
Table 4. Performance comparison between MiniOneRec and MiniOneRec* on the Industrial dataset. MiniOneRec* is obtained by replacing the original SFT-based backbone in MiniOneRec with TS-Rec.
Dataset Metric MiniOneRec MiniOneRec* Improvement
Industrial HR@3 0.0991 0.1118 12.81%
HR@5 0.1160 0.1295 11.63%
HR@10 0.1443 0.1568 8.66%
NDCG@3 0.0860 0.0995 15.68%
NDCG@5 0.0928 0.1068 15.08%
NDCG@10 0.1020 0.1156 13.44%
To evaluate whether the fine-grained semantic grounding of TS-Rec can facilitate more effective RL optimization, we integrate it into MiniOneRec (Kong et al., 2025b), a state-of-the-art open-source framework for scaling generative recommendation. Specifically, we replace MiniOneRec’s original SFT backbone with TS-Rec (denoted as MiniOneRec*) while keeping the subsequent RL (e.g., GRPO) configurations identical. The results on the Industrial dataset are summarized in Table 4, yielding several critical insights: (1) Substantial Performance Elevation across Ranking Metrics. MiniOneRec* consistently outperforms the original MiniOneRec with significant margins. Notably, we observe a 15.68% improvement in NDCG@3 and a 15.08% boost in NDCG@5. Since NDCG is highly sensitive to the exact position of items in the generated list, these gains suggest that the semantic priors injected by TS-Rec allow the RL agent to distinguish between subtly different user preferences more accurately during the exploration phase. (2) TS-Rec as a High-Quality Policy Prior. The superiority of MiniOneRec* demonstrates that TS-Rec provides a much stronger ”starting point” for RL. Unlike random initialization which forces RL to learn both token semantics and user preferences simultaneously, SA-Init and TS-Align decouple these challenges. By grounding the SID tokens in the LLM’s language space beforehand, TS-Rec reduces the search space for RL, allowing it to focus on maximizing recommendation rewards rather than struggling to comprehend the basic identity of the items. (3) Enhanced Exploration Efficiency. We observe that the relative gains in NDCG metrics are generally higher than those in HitRatio (e.g., +15.68% vs. +12.81% for @3). This indicates that TS-Rec helps the model not just ”hit” the right item, but rank it more precisely. In a generative setting, this is likely because the fine-grained token-level alignment prevents the model from generating ”semantically hallucinated” SID sequences, ensuring that the RL-tuned policy stays within a valid and meaningful item representation manifold.
In summary, the transition from SFT to RL is not a disjointed process; rather, the quality of semantic integration during SFT determines the ultimate performance ceiling of the RL phase. TS-Rec effectively raises this ceiling by creating a semantically coherent representation space, serving as an indispensable backbone for LLM-based Recommendation models.
5.5. The effect of TS-Rec on SID Comprehension (RQ4)
To evaluate whether TS‑Rec can effectively capture and utilize fine‑grained SID semantics, we design two types of generative retrieval tasks:
  • Title2SID task: Given the textual title of an item, predict its corresponding SID token sequence.
  • SID2Title task: Given the SID token sequence of an item, predict its corresponding textual title.
Figure 5. Accuracy of generative retrieval tasks on the Industrial and Office datasets. ACC@1 measures Title2SID accuracy, and ACC@2 measures SID2Title accuracy.
Both tasks rely on beam search decoding with a width of 5. We use ACC1 to denote the accuracy of Title2SID predictions, and ACC2 to measure SID2Title accuracy. We compare TS-Rec with a baseline that uses random initialization combined with item‑level SFT, evaluating on the Industrial and Office datasets. Results are summarized in Figure 5. Our key observations are as follows: (1) TS-Rec achieves substantial improvements over item-level SFT in both tasks. On Industrial, ACC1 increases from 0.3834 to 0.8469 (+120.97%), and ACC2 from 0.1599 to 0.3811 (+138.34%). On Office, ACC1 rises from 0.4783 to 0.6590 (+37.80%), and ACC2 from 0.1547 to 0.3356 (+116.79%). These improvements indicate that TS‑Rec retains and leverages SID‑level semantic information far more effectively than item‑level SFT. (2) Broader implications for recommendation. Accurately mapping between textual descriptions and structured SIDs is crucial for LLM-based recommendation systems, particularly in cold‑start scenarios, cross‑domain item transfer, and semantic search. The ability of TS-Rec to excel in generative retrieval tasks suggests that it could power recommendation models capable of seamlessly bridging natural language queries and structured semantic identifiers, thereby improving accessibility and retrieval diversity.
6. Conclusion and Future Work
In this work, we introduced TS-Rec, a novel framework for enhancing LLM-based generative recommendation through fine-grained semantic integration. By addressing two fundamental limitations of existing SID modeling, namely semantically meaningless initialization and coarse-grained alignment, TS-Rec incorporates Semantic-Aware Embedding Initialization (SA-Init) to endow SID tokens with meaningful priors and Token-level Semantic Alignment (TS-Align) to refine the alignment between SID space and the LLM’s language space. Extensive experiments on two real-world benchmarks demonstrate that TS-Rec consistently outperforms both traditional and generative baselines across all standard metrics, and further serves as a stronger backbone within SFT-then-RL pipelines. Moreover, our ablation analyses validate the complementary contributions of SA-Init and TS-Align, while in-depth studies of initialization depth reveal consistent gains in convergence speed and ranking accuracy. TS-Rec also substantially improves the LLM’s comprehension of SIDs, enabling precise bidirectional mapping between structured identifiers and natural language descriptions. This ability is critical to cold-start scenarios.
In the future, we intend to explore more sophisticated mechanisms for aligning the SID space with the language space. While our current approach integrates fine-grained semantics via static initialization and SFT-based tasks, the emerging SFT-then-RL paradigm (Kong et al., 2025b; Tan et al., 2025; You et al., 2025; Liu et al., 2025) offers a promising avenue for further optimization. We plan to investigate how token-level semantic rewards can be explicitly incorporated into the reinforcement learning phase to penalize semantically inconsistent generations and further refine the model’s preference modeling capabilities. Additionally, extending TS-Rec to multi-modal recommendation remains an open and promising direction.
Appendix A Appendix
A.1. Prompt
Prompt for Semantic Extraction # Role: You are a Senior Product Content Analyst and Catalog Specialist with expertise in {category} product classification and summarization. # Task: Analyze a provided list of diverse product items and synthesize their shared characteristics into a single, cohesive ”Product Family Description,” then produce a related keyword list. ##Input Format: Each item provides: Title: product name Description: detailed text about features, functions, and specifications Brand: brand name Categories: classification field(s) ## Core Objective: Identify Commonalities: Determine the underlying functional, design, usage, or performance traits that ALL listed items share. Filter Fluff: Exclude brand-specific marketing slogans, proprietary technology names, or promotional phrases (e.g., ”award-winning,” ”exclusive edition”). Abstract Specifications: Replace exact measurements, quantities, or configurations with generalized terms (e.g., ”various sizes,” ”multiple capacity options,” ”bulk packaging”). Synthesize: Produce a high-level overview that describes the product family as a category, rather than a list of individual items. Extract Keywords: Identify 10-20 domain-relevant, semantically meaningful keywords derived from the description. ## Style & Tone: Professional, neutral, and domain-agnostic. Objective third-person, suitable for product catalogs and technical documentation. One paragraph of 3–5 sentences that flows naturally. Clear, concise, and focused on function, durability, applicability, or end-user benefits. ## Constraints: No brand names in the output. No bullet points; provide a cohesive, continuous paragraph. Ensure the description is broad enough to encompass all items in the list but specific enough to convey meaningful shared traits. Avoid overly promotional adjectives; emphasize factual, verifiable characteristics. # Input: Items: {items}
A.2. Fine-tuning Task for Recommendation
A.2.1. Sequential Item Prediction
Instruction ($x$): The user has interacted with items ¡a_251¿¡b_198¿¡c_114¿, …, ¡a_251¿¡b_198¿¡c_132¿ in chronological order. Can you predict the next possible item that the user may expect? Response ($y$): ¡a_251¿¡b_198¿¡c_238¿
A.2.2. Explicit SID-Language Alignment
Instruction ($x$): Which item has the title: 6061 Aluminum Round Rod, Unpolished (Mill) Finish, Extruded, T6511 Temper, ASTM B221, 2D̈iameter, 12L̈ength? Response ($y$): ¡a_251¿¡b_198¿¡c_238¿
Instruction ($x$): What is the title of item ”¡a_251¿¡b_198¿¡c_238¿”? Response ($y$): 6061 Aluminum Round Rod, Unpolished (Mill) Finish, Extruded, T6511 Temper, ASTM B221, 2D̈iameter, 12L̈ength
A.2.3. Asymmetric Item Prediction-1
Instruction ($x$): The user has sequentially interacted with items history. Can you recommend the next item for him? Tell me the title of the item. Response ($y$): 6061 Aluminum Round Rod, Unpolished (Mill) Finish, Extruded, T6511 Temper, ASTM B221, 2D̈iameter, 12L̈ength
A.2.4. Asymmetric Item Prediction-2
Instruction ($x$): Please review the user’s historical interactions: ¡a_251¿¡b_198¿¡c_114¿, …, ¡a_251¿¡b_198¿¡c_132¿, and describe what kind of item he still needs. Response ($y$): The 6061 aluminum round rod with T6511 temper has an unpolished (mill) surface, …, while weldability characterizes the ability to be welded.

### v2

@@
-Fine-grained Semantics Integration for Large Language Model-based Recommendation
+arXiv:2602.22632v2[cs.IR] 28 Feb 2026
 Abstract.
@@
 Building on this paradigm, LLM-based generative recommendation further incorporates large language models to leverage their rich world knowledge, enhancing the recommender’s understanding of items and thereby improving recommendation performance(Liu et al., 2025; Han et al., 2025; Team, 2025; Zhang et al., 2023).
-In a typical LLM-based generative recommendation pipeline (Liu et al., 2025; Kong et al., 2025b; You et al., 2025), the training procedure can be broadly divided into three stages: (1) item tokenization as semantic identifiers (SIDs)(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023): each item is first represented as a sequence of $N$ fixed-length tokens (i.e., SIDs) that lie outside the original vocabulary of the LLM (e.g., three tokens like ¡a_236¿ ¡b_231¿ ¡c_226¿), commonly obtained using vector quantization methods such as RQ-VAE(Zeghidour et al., 2021); (2) alignment(Zheng et al., 2024; Wang et al., 2024) between these out-of-vocabulary (OOV) item tokens and the LLM backbone through supervised fine-tuning (SFT), enabling the model to understand these tokens for recommendation; and (3) preference learning (e.g., reinforcement learning (RL)(Tan et al., 2025; Kong et al., 2025b; Team, 2025; Zhou et al., 2025; Kong et al., 2025a; You et al., 2025) or direct preference optimization (DPO) (Chen et al., 2024; Bai et al., 2024; Rafailov et al., 2023; Ye et al., 2025)) to further enhance the user preferences modeling capabilities, thereby pushing the upper bound of recommendation quality.
+In a typical LLM-based generative recommendation pipeline (Liu et al., 2025; Kong et al., 2025b; You et al., 2025), the training procedure can be broadly divided into three stages: (1) item tokenization as semantic identifiers (SIDs)(Rajput et al., 2023; Wang et al., 2024; Hou et al., 2023): each item is first represented as a sequence of NN fixed-length tokens (i.e., SIDs) that lie outside the original vocabulary of the LLM (e.g., three tokens like ¡a_236¿ ¡b_231¿ ¡c_226¿), commonly obtained using vector quantization methods such as RQ-VAE(Zeghidour et al., 2021); (2) alignment(Zheng et al., 2024; Wang et al., 2024) between these out-of-vocabulary (OOV) item tokens and the LLM backbone through supervised fine-tuning (SFT), enabling the model to understand these tokens for recommendation; and (3) preference learning (e.g., reinforcement learning (RL)(Tan et al., 2025; Kong et al., 2025b; Team, 2025; Zhou et al., 2025; Kong et al., 2025a; You et al., 2025) or direct preference optimization (DPO) (Chen et al., 2024; Bai et al., 2024; Rafailov et al., 2023; Ye et al., 2025)) to further enhance the user preferences modeling capabilities, thereby pushing the upper bound of recommendation quality.
 As the first two steps, item tokenization and alignment with LLMs directly determine how the LLM understands items and thus largely affect recommendation performance, playing vital roles in LLM-based generative recommendation.
@@
 Despite these advancements, traditional SID-based generative methods only encode item semantics into the random initialized SID embeddings, while largely overlooking the potential of exploiting the inherent semantic capabilities of the generative backbone.
-2.2. LLM-based Generative Recommendation
+$2.$ LLM-based Generative Recommendation
 Another line of work focuses on utilizing modern LLMs (Yang et al., 2024; Team, 2024) as the recommendation backbone to harness their vast semantic understanding and world knowledge (Bao et al., 2023; Liao et al., 2024; Karra and Tulabandhula, 2024; Liu et al., 2025; Team, 2025; Kong et al., 2024).
@@
 The continuous embedding is subsequently discretized using the RQ-KMeans algorithm (Luo et al., 2025).
-Let $\tilde{\mathbf{M}}=[\mathbf{x}_{1};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ be the matrix stacking the embeddings of all $N$ items.
+Let $\tilde{\mathbf{M}}=[\mathbf{x}_{1};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ be the matrix stacking the embeddings of all NN items.
 We initialize $\mathbf{R}^{(1)}=\tilde{\mathbf{M}}$.
@@
 Here, $\mu$ and $\Sigma$ denote the mean vector and covariance matrix of the LLM’s pre-trained embedding space.
-3.3. Supervised Fine Tuning for LLM-based Recommenders
+$3.$ Supervised Fine Tuning for LLM-based Recommenders
 In supervised fine-tuning (SFT), the semantic understanding and generative capabilities of large language models (LLMs) are leveraged to enhance sequential recommendation.
@@
     (6) $\mathcal{I}_{s}=\{i\in\mathcal{I}\mid s\in e_{i}\},$
-    where $e_{i}$ represents the SID tuple of item $i$. This step groups items based on shared structural codes, serving as a proxy for shared semantics.
+    where $e_{i}$ represents the SID tuple of item ii. This step groups items based on shared structural codes, serving as a proxy for shared semantics.
   • Semantic Extraction.
@@
 TS-Rec effectively raises this ceiling by creating a semantically coherent representation space, serving as an indispensable backbone for LLM-based Recommendation models.
-5.5. The effect of TS-Rec on SID Comprehension (RQ4)
+$5.$ The effect of TS-Rec on SID Comprehension (RQ4)
 To evaluate whether TS‑Rec can effectively capture and utilize fine‑grained SID semantics, we design two types of generative retrieval tasks: